OpenStudy (goformit100):

The length x of a rectangle is decreasing at the rate of 5 cm/minute and the width y is increasing at the rate of 4 cm/minute. When x = 8cm and y = 6cm, find the rates of change of (a) the perimeter , and (b) the area of the rectangle.

OpenStudy (goformit100):

@terenzreignz

OpenStudy (goformit100):

@genius12

OpenStudy (anonymous):

\[\frac{ dx }{ dt }=-5 \ cm/\min\]\[\frac{ dy }{ dt }=4 \ cm/\min\]\[A = xy \ and \ P =2x+2y\]\[\frac{ dA }{ dt }=y \frac{ dx }{ dt }+x \frac{ dy }{ dt }\]\[\frac{ dP }{ dt }=2\frac{ dx }{ dt }+2\frac{ dy }{ dt }\]Now just plug everything in and solve. @goformit100 Also it would be nice if you actually tried to do some of these yourself rather than posting questions for each one. That's only way for you to learn.

OpenStudy (goformit100):

ThanKyou Sir